OpenStudy (bamagirl04):

Helpp ill fan and medal!! m-12=16

OpenStudy (itsbribro):

you can add 16 and 12 together and get your answer :)

OpenStudy (anonymous):

m-12=16 m-12+12=16+12 m=28